OpenStudy (anonymous):

Suppose that a + b = 3 and a^2 + b^2 = 7. Then a^4 + b^4 is equal to

OpenStudy (anonymous):

\[(a^2+b^2)^2=7^2\]

OpenStudy (anonymous):

\[a^4+2a^2b^2+b^4=49\]

OpenStudy (anonymous):

is (a^2 + b^2)^2 = a^4 + b^4 ?

OpenStudy (anonymous):

\[a^4+b^4+2a^2b^2=49 => a^4+b^4=49-2a^2b^2\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it is what i said

OpenStudy (anonymous):

\[(a^2+b^2)^2=a^4+2a^2b^2+b^4\]

OpenStudy (anonymous):

i see ..

OpenStudy (anonymous):

\[(a+b)^2=3^2 => a^2+2ab+b^2=9\]

OpenStudy (anonymous):

\[a^2+b^2=9-2ab\] But what does \[a^2+b^2=?\]

OpenStudy (anonymous):

7 right?

OpenStudy (anonymous):

so 9-2ab=7

OpenStudy (anonymous):

solve for ab

OpenStudy (anonymous):

\[-2ab=-2 => ab=1\]

OpenStudy (anonymous):

so if ab=1 then \[(ab)^2=?\]

OpenStudy (anonymous):

Remember we have this \[a^4+b^4=49-2a^2b^2\]

OpenStudy (anonymous):

we know \[(ab)^2=a^2b^2\]

OpenStudy (anonymous):

so if \[ab=1 \text{ what is } (ab)^2=?\] @cuty_shai2000 ?

OpenStudy (anonymous):

a^4 + b^4 = 47?

OpenStudy (anonymous):

49-2 = 47? hmm

OpenStudy (anonymous):

yep yep :)

OpenStudy (anonymous):

now i know, thanks so much.

OpenStudy (anonymous):

np